OpenStudy (anonymous):

I have to apply the Green's theorem to a triangle. I've got the integral of e^xcosy dx-e^xsiny dy. A=(ln2,0), C=(0,1) and B(-ln2,0). Can someone help me please? Thank you.

OpenStudy (reemii):

\(F(x,y)=(e^x\cos y, -e^x\sin y)=(P,Q) \) Green: \[ \text{your integral} = \int\int_\text{surface of triangle} (\frac{\partial Q}{\partial y}-\frac{\partial P}{\partial x}) dxdy \] You use Fubuni to write correctly \(\int\int_\text{surface of triangle}\), that is, you can choose the order of integration. Here it is best to integrate on two different triangles: the two that you get when you cut along the \(y\)-axis.

OpenStudy (reemii):

your integral = \[\int_0^{\ln 2}(\int_0^{1-x} H(x,y) dy)dx + \int_{-\ln 2}^0(\int_0^{x+1} H(x,y) dy)dx,\] where \(H(x,y)=(∂Q/∂y−∂P/∂x)\)

OpenStudy (reemii):

sorry I made a mistake with the integrals..

OpenStudy (reemii):

first integral: \[ \int_0^{\ln 2}(\int_a^b H(x,y) \,dy)dx \] obviously, \(a=0\), but \(b=?\) 

OpenStudy (reemii):

okay this is right: \[ \int_{\ln2}^0 (\int_0^{b} H(x,y) \:dy)dx \]where \(b\) is y_max. the equation of the line is \(y=(1/(-\ln2))x+1\). That is \(b\).

OpenStudy (anonymous):

@reemii is it that? is it not line integral?

OpenStudy (reemii):

The statement is an integral over a path in \(\mathbb{R}^2\), but the Green formula enables to write an integral on a surface (the delimited by the closed path).

OpenStudy (anonymous):

hold on, look at the function, it's .....dx - ..... dy, is it not the form of line integral? Please, don't confuse me. I need that part for my final, too.

OpenStudy (reemii):

I'm sure. what he wrote : "e^xcosy dx-e^xsiny dy on a triangle ABC" is to be taken as a notation. I think the word for that topic is "flow integral". Stated differently, the starting point is: \(\int_\text{triangle ABC} \text{flow} F \cdot \mathrm ds\). do you recognize this?

OpenStudy (anonymous):

the question is unclear about the triangle. He disappeared , we cannot confirm the question but to me, it looks like line integral with parametric equation from A to B and then from B to C . Moreover, the form of function is not dA =dx dy but ....dx - ....dy, separately.

OpenStudy (reemii):

this is a notation. the fact that i'm pretty sure i'm right is beecause he mentions Green, this means the path is closed. The notation "e^xcosy dx-e^xsiny dy" means that he has to compute the flow of the vector field (e^xcosy , -e^xsiny) along the path described be the triangle.

OpenStudy (anonymous):

thanks for being here, the Asker,

OpenStudy (reemii):

oh, sorry i'm not aware of the vocabulary. It seems the starting point was indeed : line integral of a vector field.

OpenStudy (anonymous):

okk guys. My question is: integralo f C e^xcosydx-e^xsinydy where Cis the broken line from A=(ln2,0) to D=(0,1) and then from D to B=(-ln2,0). Hint: Apply Green's theorem to the integral around the closed curve ADBA.

OpenStudy (anonymous):

ok, I am waiting for other's solution, no idea

OpenStudy (anonymous):

The answer that I need to get is -3/2 . But something that I am doing is wrong cause I 'm finding other answers.

OpenStudy (reemii):

ok maybe don't cut the triangle. Integrate like this: \(\int (\int ... dx )dy\).